Suppose you physically simulate the random process of rolling a single die. (a) After 10 rolls of the die, you observe a one 2 times. What proportion of the rolls resulted in a one? (b) After 20 rolls of the die, you observe a one 6 times. What proportion of the rolls resulted in a one? (a) The proportion was Simplify your answer. Do not round.)
Answer #1

a)proportion of the rolls resulted in a "one" =2/10 =1/5

b)proportion of the rolls resulted in a "one" =6/20 =3/10